OpenStudy (anonymous):

indicate in standard form the equation of the line passing through the given points (4,1) (5,2)

OpenStudy (y1066380):

1. calculate the slope of 2 points 2. Use the equation y-y1=m(x-x1)

OpenStudy (anonymous):

would it be x-6y=-5